# PathSeeker: Exploring LLM Security Vulnerabilities with a Reinforcement Learning-Based Jailbreak Approach

### v1

@@ Abstract.
In recent years, Large Language Models (LLMs) have gained widespread use, accompanied by increasing concerns over their security. Traditional jailbreak attacks rely on internal model details or have limitations when exploring the unsafe behavior of the victim model, limiting their generalizability. In this paper, we introduce PathSeeker, a novel black-box jailbreak method inspired by the concept of escaping a security maze. This work is inspired by the game of rats escaping a maze. We think that each LLM has its unique “security maze”, and attackers attempt to find the exit learning from the received feedback and their accumulated experience to compromise the target LLM’s security defences. Our approach leverages multi-agent reinforcement learning, where smaller models collaborate to guide the main LLM in performing mutation operations to achieve the attack objectives. By progressively modifying inputs based on the model’s feedback, our system induces richer, harmful responses. During our manual attempts to perform jailbreak attacks, we found that the vocabulary of the response of the target model gradually became richer and eventually produced harmful responses. Based on the observation, we also introduce a reward mechanism that exploits the expansion of vocabulary richness in LLM responses to weaken security constraints. Our method outperforms five state-of-the-art attack techniques when tested across 13 commercial and open-source LLMs, achieving high attack success rates, especially in strongly aligned commercial models like GPT-4o-mini, Claude-3.5, and GLM-4-air with strong safety alignment. This study aims to improve the understanding of LLM security vulnerabilities and we hope that this sturdy can contribute to the development of more robust defenses.
@@ 1. Introduction
Large Language Models (LLMs) demonstrate the potential for general artificial intelligence (Bubeck et al., 2023). However, LLMs also face security threats (Yao et al., 2024b). Currently, the primary approach to addressing LLM security threats is through safety alignment techniques (Bai et al., 2022, Zhao et al., 2024, Song et al., 2024), ensuring that the output of LLMs aligns with human ethics and values. To test the security of LLMs, various jailbreak techniques (Jin et al., 2024a) have been proposed, such as GCG (Zou et al., 2023), CodeChameleon (Lv et al., 2024), GPTFuzzer (Yu et al., 2023), RL-JACK (Chen et al., 2024) and RLTA (Wang et al., 2024). These techniques can be categorized into white-box attacks, which rely on internal model information, and black-box attacks, which do not. Black-box attacks, in particular, do not require access to the model’s internal details, making them applicable to a wide range of models. Most black-box attack algorithms depend on LLMs analyzing the output of the target model to improve the next attack. GPTFuzzer (Yu et al., 2023), RL-JACK (Chen et al., 2024) and RLTA (Wang et al., 2024) are most close to our work while we are definitely different, and our solution is more general. GPTFuzzer uses a fine-tuned small model in the attack loop but does not leverage the experience gained from each attack attempt. RL-JACK and RLTA aim to improve the effectiveness of black-box attacks by incorporating intermediate feedback during the attack process. However, both approaches rely on harmful reference answers to obtain feedback. They use proxy models to generate answers to the harmful questions, and these answers are then used to guide the target model by computing BLEU scores or cosine similarity. There are two main limitations: (1) The target model will produce answers similar to the proxy model, which limits the ability to test the target model’s creativity. In an attack scenario, creativity refers to how the target model generates harmful outcomes; (2) When the unsafe behaviors of the target and proxy models have little overlap, it becomes difficult to breach the target model since the attack assumes the target model will produce answers similar to the proxy model. In the following paragraphs, we present a motivational example to demonstrate these two limitations.
@@
To vividly illustrate the advantages of our method compared to other reinforcement learning-based approaches, we use this game as an example. Solutions based on reference answers from proxy models (Chen et al., 2024, Wang et al., 2024) are akin to constructing a simulated maze that represents the real security maze of the target models. When the simulated maze closely resembles the real one, the rat can find a similar route to the exit using its experience from the simulated maze. In contrast, when the two mazes are significantly different or the real maze is far more complex due to strong safety alignment, the experience from the simulated maze offers little value in helping the rat escape the real one. In essence, the difference between our approach and others is like between the supervised learning and unsupervised learning, indicating by if one approach needs the answer to the harmful question under attack.
@@ 3.2. Experimental Setup
We consider 5 SOTA jailbreak attack methods which are open source. The concurrent works, RL-JACK (Chen et al., 2024) and RLTA (Wang et al., 2024), are not open source at the moment and we do not inlcude them.

### v2

@@ Abstract.
In recent years, Large Language Models (LLMs) have gained widespread use, raising concerns about their security. Traditional jailbreak attacks, which often rely on the model internal information or have limitations when exploring the unsafe behavior of the victim model, limiting their reducing their general applicability. In this paper, we introduce PathSeeker, a novel black-box jailbreak method, which is inspired by the game of rats escaping a maze. We think that each LLM has its unique “security maze”, and attackers attempt to find the exit learning from the received feedback and their accumulated experience to compromise the target LLM’s security defences. Our approach leverages multi-agent reinforcement learning, where smaller models collaborate to guide the main LLM in performing mutation operations to achieve the attack objectives. By progressively modifying inputs based on the model’s feedback, our system induces richer, harmful responses. During our manual attempts to perform jailbreak attacks, we found that the vocabulary of the response of the target model gradually became richer and eventually produced harmful responses. Based on the observation, we also introduce a reward mechanism that exploits the expansion of vocabulary richness in LLM responses to weaken security constraints. Our method outperforms five state-of-the-art attack techniques when tested across 13 commercial and open-source LLMs, achieving high attack success rates, especially in strongly aligned commercial models like GPT-4o-mini, Claude-3.5, and GLM-4-air with strong safety alignment. This study aims to improve the understanding of LLM security vulnerabilities and we hope that this sturdy can contribute to the development of more robust defenses.
@@ 1. Introduction
Large Language Models (LLMs) demonstrate the potential for general artificial intelligence (Bubeck et al., 2023). However, LLMs also face security threats (Yao et al., 2024b). Currently, the primary approach to addressing LLM security threats is through safety alignment techniques (Bai et al., 2022, Zhao et al., 2024, Song et al., 2024), ensuring that the output of LLMs aligns with human ethics and values. To test the security of LLMs, various jailbreak techniques (Jin et al., 2024a) have been proposed, such as GCG (Zou et al., 2023), CodeChameleon (Lv et al., 2024), GPTFuzzer (Yu et al., 2023), RLbreaker (Chen et al., 2024a), RL-JACK (Chen et al., 2024b) and RLTA (Wang et al., 2024). These techniques can be categorized into white-box attacks, which rely on internal model information, and black-box attacks, which do not. Black-box attacks, in particular, do not require access to the model’s internal details, making them applicable to a wide range of models. Most black-box attack algorithms depend on LLMs analyzing the output of the target model to improve the next attack. GPTFuzzer (Yu et al., 2023), RLbreaker (Chen et al., 2024a), RL-JACK (Chen et al., 2024b) and RLTA (Wang et al., 2024) are most close to our work while we are definitely different, and our solution is more general. GPTFuzzer uses a fine-tuned small model in the attack loop but does not leverage the experience gained from each attack attempt. RLbreaker, RL-JACK and RLTA aim to improve the effectiveness of black-box attacks by incorporating intermediate feedback during the attack process. However, both approaches rely on harmful reference answers to obtain feedback. They use proxy models to generate answers to the harmful questions, and these answers are then used to guide the target model by computing BLEU scores or cosine similarity. There are two main limitations: (1) The target model will produce answers similar to the proxy model, which limits the ability to test the target model’s creativity. In an attack scenario, creativity refers to how the target model generates harmful outcomes; (2) When the unsafe behaviors of the target and proxy models have little overlap, it becomes difficult to breach the target model since the attack assumes the target model will produce answers similar to the proxy model. In the following paragraphs, we present a motivational example to demonstrate these two limitations.
@@
To vividly illustrate the advantages of our method compared to other reinforcement learning-based approaches, we use this game as an example. Solutions based on reference answers from proxy models (Chen et al., 2024b, Wang et al., 2024) are akin to constructing a simulated maze that represents the real security maze of the target models. When the simulated maze closely resembles the real one, the rat can find a similar route to the exit using its experience from the simulated maze. In contrast, when the two mazes are significantly different or the real maze is far more complex due to strong safety alignment, the experience from the simulated maze offers little value in helping the rat escape the real one. In essence, the difference between our approach and others is like between the supervised learning and unsupervised learning, indicating by if one approach needs the answer to the harmful question under attack.
@@ 3.2. Experimental Setup
We consider 5 SOTA jailbreak attack methods which are open source. The concurrent works, RL-JACK (Chen et al., 2024b) and RLTA (Wang et al., 2024), are not open source at the moment and we do not inlcude them.